Using the table of average bond enthalpies, what is the ∆H for the following gas phase...

  1. Using the table of average bond enthalpies, what is the ∆H for the following gas phase reaction (in kJ/mol)?

    ethene reacting with hydrobromic acid to produce bromoethane

    Bond Ave. Bond Enthalpy (kJ/mol)
    C – C 348
    C = C 614
    C ≡ C 839
    C – H 413
    C – Br 276
    H – Br 366
    H – H 436
    a

    - 304

    b

    + 304

    c

    - 2134

    d

    - 57

    e

    + 3292
Answer #1

HL - H I- n- - I Here the gas phase reaction equ- ation can be given as # H 1 c = c H (9) + e H-Bro (9) la (ethene). # it → HAgain . E (BE) product = (5x C-H bond enthalpy) + (1xc-c bond enthalpy) + (1x c-Br bond enthalpy) = [(5x 413) + 348 + 276] kJThus, SH = (2632 - 2689) kJ/m01 = -57 So, the correct option would be 2. -57) An